# Neutrinos in colliding neutron stars and black holes

### v1

@@ III.1.2 Pair processes
and nucleon-nucleon Bremsstrahlung. Here, the neutrinos can be of any flavor, though the exact reaction rate varies with neutrino flavor. The forward reaction depends mostly of the availability of electron-positron pairs, and the reaction rate will thus increase steeply with temperature (the energy emission rate scales as $T^{9}$). For the typical conditions encountered in mergers, pair production of $\nu_{e}\bar{\nu}_{e}$ is typically slower than emission through charged current reactions. For production of muon and tau neutrinos, given the low rate of charged current reactions, this will often be the dominant process. As a result, heavy-lepton neutrinos decouple from the fluid at higher densities and temperature than electron type neutrinos within post-merger neutron stars (see Fig. 3), and are rarely absorbed within post-merger accretion disks.
@@ III.1.3 Scattering
All species of neutrinos additionally interact with the matter through scattering onto neutrons, protons, electrons and nuclei. The scattering opacity for collisions on neutrons and protons can be approximated as [33, 35, 36]
@@ III.2 Neutrino Flavor Conversion
In addition to their interactions with matter, the distribution of neutrinos in neutron star mergers may be impacted by flavor conversion. The fact that the mass eigenstates of neutrinos are a superposition of all three neutrino flavors is well known, and results in vacuum oscillations as well as flavor oscillations within slowly varying matter density profiles (MSW mechanism) that were, for example, crucial to understand the neutrino signal coming from the Sun [39, 40, 41, 42]. Flavor oscillations over length scales much larger than the size of the remnant neutron stars can modify the luminosity of neutrinos of different flavor as observed from Earth. Unfortunately, for the neutrinos emitted from the neutron star remnant and disk at least, such direct observations are unlikely to be possible lacking an extremely lucky close-by event (i.e. within our galaxy or local group). For example, water detectors such as Hyper-Kamiokande may need to monitor thousands of merger to detect a single neutrino [43].
@@
Other drivers of neutrino flavor conversion can be more important very close to the merger remnant, in regions where neutrino-matter interactions can still significantly impact the properties of matter outflows. In the presence of high neutrino and/or matter densities, modifications to the Hamiltonian describing the evolution of neutrinos can cause flavor conversions on shorter length scales. The neutrino-matter resonance (NMR) [44, 45] occurs when the matter potential and neutrino self-interaction potential intersects, and can be active within a few neutron star radii of a post-merger remnant, though the amount of flavor conversion from the NMR may be insufficient to significantly impact observables from neutron star mergers [46]. The fast-flavor instability (FFI) [47, 48, 49, 50] is purely due to collective neutrino interactions. Neutrinos are expected to be unstable to the FFI as soon as a crossing in the net electron neutrino flux occurs (i.e. at a given point, the flux of neutrinos is larger than the flux of antineutrinos in some directions, but smaller in other directions) [51, 52, 53]. The FFI has been of particular interest in recent years because neutrinos directly outside of a neutron star remnant and close to the disk can be unstable to the FFI, and flavor conversions close to the remnant are most likely to modify the relative fluxes of $\nu_{e}$ and $\bar{\nu}_{e}$ in ways that would impact absorption in the matter outflows, and thus the neutron-richness of these outflows. Existing global simulations of neutron star mergers at best include very approximate treatments of the FFI [54, 55, 56]. These results indicate that the FFI may modify predictions for r-process nucleosynthesis, but by how much remains an open question. Finally, the collisional instability [57] is precipitated in denser regions of the remnant by unequal neutrino and antineutrino scattering rates, and may also be active close to the neutrino decoupling regions. It remains unclear whether it can practically impact the outcome of neutron star mergers. In the context of core-collapse supernovae, recent studies have found that the impact of collisional instabilities may be limited [58, 59].
@@ III.3.1 Cooling of the remnant
Emission from the accretion disk may remain as high as $10^{52-53}\,{\rm erg/s}$ for $O(100\,{\rm ms})$, the timescale for viscous spreading of the disk [28, 66, 67]. At this luminosity and for typical accretion rates in post-merger remnants, the change in the binding energy of the matter as it accretes on the central object is slightly larger than the energy lost to neutrinos, so that part of that energy serves to heat the accretion disk. An efficiently cooled accretion disk would become very thin (height over radius ratio $H/r\ll 1$), due to the lack of vertical pressure support. This is seen for example in thin disks around supermassive black holes, which can be efficiently cooled by photon emission. In post-merger remnants, neutrino cooling does help to reduce the height of the disk, but only down to $H/r\sim 0.2$. Neutrino cooling has an important effect on the evolution of the disk, but is not sufficient to efficiently radiate the energy generate by the accretion process. After $\sim 0.1\,{\rm s}$, as the accretion rate decreases, neutrino cooling becomes less efficient. Indeed, the main cooling reactions $n+e^{+}\rightarrow p+\bar{\nu}_{e}$ and $p+e^{-}\rightarrow n+\nu_{e}$ have luminosities scaling as $(n_{n}n_{e^{+}})$ and $(n_{p}n_{e^{-}})$, respectively, while the thermal energy of the disk approximately scales linearly with its density. As soon as the disk density decreases, neutrino cooling becomes inefficient and the accretion disk puffs up to $H/r\sim 1$ [68, 69]. Viscous heating and $\alpha$-particle formation release energy, with no cooling mechanism to compensate. Over seconds timescale, this leads to the ejection of a significant fraction of the disk mass ($\sim 5\%-20\%$[70].
@@ III.3.2 Changes in the composition of the fluid
The second important effect of neutrinos in post-merger remnants, and the one with the most direct impact on observables, is to drive changes in the composition of the post-merger remnant, i.e. its neutron-richness. Before merger, most of the matter has $Y_{e}\lesssim 0.1$. As soon as the density of the matter decreases, or the temperature increases, it becomes energetically favorable for $Y_{e}$ to increase. Charged current reaction transforming neutrons into protons are favored over their inverse reactions. In the cold matter ejected during the tidal disruption of a neutron star, neutrino emission and absorption are rare. Accordingly, $Y_{e}$ does not change much, and the matter remains very neutron rich. In the accretion disk and in the hot outflows ejected from the disk and neutron star remnant, on the other hand, neutrino-matter interactions quickly lead to an increase in $Y_{e}$ [72]. Fig. 4 illustrates the large difference in composition between a remnant neutron star (very neutron rich), an accretion disk (less neutron rich), and hot matter ouflows ($Y_{e}>0.3$).
Figure 4: Vertical slice through a post-merger neutron star remnant $10\,{\rm ms}$ after merger (from the same simulation as Fig. 3). The dotted white lines show density contours of $10^{11,12,13}\,{\rm g/cm^{3}}$; the solid black line marks the boundary between bound and ejected matter. Arrows show the velocity of the fluid, and the color scale shows the composition of the remnant. We see the highly neutron-rich core at the center, the moderately neutron-rich disk in the equatorial plane, and higher $Y_{e}$ outflows in the polar regions. Figure reproduced from [38]
In the accretion disk, charged current reactions are initially fast enough for the matter to reach its equilibrium composition of $Y_{e}\sim(0.1-0.2)$ within a few milliseconds. At later times, the composition of the disk evolves because the equilibrium composition varies with the density and temperature of the matter, increasing with decreasing density. After $O(0.1\,{\rm s})$, weak interaction become too slow to maintain the equilibrium electron fracition, and the $Y_{e}$ of the disk freezes out. The expected value of $Y_{e}$ at freeze-out is $\sim 0.3$ [68, 69]. This is a particularly important process when considering r-process nucleosynthesis and kilonova emission. As mentioned earlier, $\sim 5\%-20\%$ of the initial accretion disk is unbound during the ADAF phase, and that matter will be ejected with an electron fraction close to the freeze-out value.
@@ Acknowledgements.
We thank Rodrigo Fernandez, Irene Tamborra, and Somdutta Ghosh for discussions and comments provided on drafts of this chapter. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. FF gratefully acknowledges support from the Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-AC02-05CH11231 and DE-SC0025023, from NASA through grant 80NSSC22K0719, and from the NSF through grant AST-2107932.

### v2

@@ III.1.2 Pair processes
and nucleon-nucleon Bremsstrahlung. Here, the neutrinos can be of any flavor, though the exact reaction rate varies with neutrino flavor. The forward reaction depends mostly of the availability of electron-positron pairs, and the reaction rate will thus increase steeply with temperature (the energy emission rate scales as $T^{9}$). For the typical conditions encountered in mergers, pair production of $\nu_{e}\bar{\nu}_{e}$ is typically slower than emission through charged current reactions. For production of muon and tau neutrinos, given the low rate of charged current reactions, this will often be the dominant process. As a result, heavy-lepton neutrinos decouple from the fluid at higher density and temperature than electron type neutrinos within post-merger neutron stars (see Fig. 3), and are rarely absorbed within post-merger accretion disks.
@@ III.1.3 Scattering
All species of neutrinos additionally interact with the matter through scattering onto neutrons, protons, electrons and nuclei. The scattering opacity (inverse mean-free path) for collisions on neutrons and protons can be approximated as [35, 37, 38]
@@ III.2 Neutrino Flavor Conversion
In addition to their interactions with matter, the distribution of neutrinos in neutron star mergers may be impacted by flavor conversion. The fact that the mass eigenstates of neutrinos are a superposition of all three neutrino flavors is well known, and results in vacuum oscillations as well as flavor oscillations within slowly varying matter density profiles (MSW mechanism) that were, for example, crucial to understand the neutrino signal coming from the Sun [41, 42, 43, 44]. Flavor oscillations over length scales much larger than the size of the remnant neutron stars can modify the luminosity of neutrinos of different flavor as observed from Earth. Unfortunately, for the neutrinos emitted from the neutron star remnant and disk at least, such direct observations are unlikely to be possible lacking an extremely lucky close-by event (i.e. within our galaxy or local group). For example, water detectors such as Hyper-Kamiokande may need to monitor thousands of mergers to detect a single neutrino [45].
@@
Other drivers of neutrino flavor conversion can be more important very close to the merger remnant, in regions where neutrino-matter interactions can still significantly impact the properties of matter outflows. In the presence of high neutrino and/or matter densities, modifications to the Hamiltonian describing the evolution of neutrinos can cause flavor conversions on shorter length scales. The neutrino-matter resonance (NMR) [46, 47] occurs when the matter potential and neutrino self-interaction potential intersects, and can be active within a few neutron star radii of a post-merger remnant, though the amount of flavor conversion from the NMR may be insufficient to significantly impact observables from neutron star mergers [48]. The fast-flavor instability (FFI) [49, 50, 51, 52] is purely due to collective neutrino interactions, i.e. it arises because of corrections to the Hamiltonian for the evolution of neutrinos that depend on the distribution function of other neutrinos. Neutrinos are expected to be unstable to the FFI as soon as a crossing in the net electron neutrino flux occurs (i.e. at a given point, the flux of neutrinos is larger than the flux of antineutrinos in some directions, but smaller in other directions) [53, 54, 55, 56]  [^3] Or rather, instability to the FFI occurs when there is a crossing in the net electron neutrino flux minus the net muon neutrino flux, but the latter is significantly smaller than the former in the merger context, where the distribution functions of $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ are fairly similar before the onset of flavor oscillations. . The result of the FFI is typically neutrino flavor conversion that erases such crossings [54, 57]. The FFI has been of particular interest in recent years because neutrinos directly outside of a neutron star remnant and close to the disk can be unstable to the FFI. Specifically, $\nu_{e}$ and $\bar{\nu}_{e}$ decouple from the fluid in different regions, and are produced in different ratios in the disk and remnant neutron star, potentially causing crossings in the net electron neutrino flux outside of the remnant. This is important because flavor conversions close to the remnant are most likely to modify the relative fluxes of $\nu_{e}$ and $\bar{\nu}_{e}$ in ways that would impact absorption in the matter outflows, and thus the neutron-richness of these outflows. Existing global simulations of neutron star mergers at best include very approximate treatments of the FFI [58, 59, 60]. These results indicate that the FFI may modify predictions for r-process nucleosynthesis, but by how much remains an open question. Finally, the collisional instability [61] is precipitated in denser regions of the remnant by unequal neutrino and antineutrino scattering rates, and may also be active close to the neutrino decoupling regions. It remains unclear whether it can practically impact the outcome of neutron star mergers. In the context of core-collapse supernovae, recent studies have found that the impact of collisional instabilities may be limited [62, 63].
@@ III.3.1 Cooling of the remnant
Emission from the accretion disk may remain as high as $10^{52-53}\,{\rm erg/s}$ for $O(100\,{\rm ms})$, the timescale for viscous spreading of the disk [30, 70, 71]. To assess the impact of neutrino cooling, it is good to compare that energy to the rate at which the accretion disk gains energy due to the increase of the binding energy of particles as they approach the compact object, i.e. $\sim G\dot{M}M_{\rm rem}/R$ (with $\dot{M}$ the accretion rate, $M_{\rm rem}$ the mass of the remnant, and $R$ the radius of the inner edge of the disk). At this luminosity and for typical accretion rates in post-merger remnants (a few percents of a solar mass per second or more in the first $0.1$ second [72]), the change in the binding energy of the matter as it accretes on the central object is slightly larger than the energy lost to neutrinos, so that part of that energy serves to heat the accretion disk. An efficiently cooled accretion disk would become very thin (height over radius ratio $H/r\ll 1$), due to the lack of vertical pressure support. This is seen for example in thin disks around supermassive black holes, which can be efficiently cooled by photon emission. In post-merger remnants, neutrino cooling does help to reduce the height of the disk, but only down to $H/r\sim 0.2$. Neutrino cooling has an important effect on the evolution of the disk, but is not sufficient to efficiently radiate the energy generated by the accretion process. After $\sim 0.1\,{\rm s}$, as the accretion rate decreases, neutrino cooling becomes less efficient. Indeed, the main cooling reactions $n+e^{+}\rightarrow p+\bar{\nu}_{e}$ and $p+e^{-}\rightarrow n+\nu_{e}$ have luminosities scaling as $(n_{n}n_{e^{+}})$ and $(n_{p}n_{e^{-}})$, respectively, while the thermal energy of the disk approximately scales linearly with its density. As soon as the disk density decreases, neutrino cooling becomes inefficient and the accretion disk puffs up to $H/r\sim 1$ [73, 74]. Viscous heating and $\alpha$-particle formation release energy, with no cooling mechanism to compensate. Over seconds timescale, this leads to the ejection of a significant fraction of the disk mass ($\sim 5\%-20\%$[75].
@@ III.3.2 Changes in the composition of the fluid
The second important effect of neutrinos in post-merger remnants, and the one with the most direct impact on observables, is to drive changes in the composition of the post-merger remnant, i.e. its neutron-richness. Before merger, most of the matter has $Y_{e}\lesssim 0.1$. As soon as the density of the matter decreases, or the temperature increases, it becomes energetically favorable for $Y_{e}$ to increase. Charged current reactions transforming neutrons into protons are favored over their inverse reactions. In the cold matter ejected during the tidal disruption of a neutron star, neutrino emission and absorption are rare. Accordingly, $Y_{e}$ does not change much, and the matter remains very neutron rich. In the accretion disk and in the hot outflows ejected from the disk and neutron star remnant, on the other hand, neutrino-matter interactions quickly lead to an increase in $Y_{e}$ [77]. Fig. 4 illustrates the large difference in composition between a remnant neutron star (very neutron rich), an accretion disk (less neutron rich), and hot matter ouflows ($Y_{e}>0.3$).
Figure 4: Vertical slice through a post-merger neutron star remnant $10\,{\rm ms}$ after merger (from the same simulation as Fig. 3). The dotted white lines show density contours of $10^{11,12,13}\,{\rm g/cm^{3}}$; the solid black line marks the boundary between bound and ejected matter. Arrows show the in-plane components of the velocity of the fluid, and the color scale shows the composition of the remnant. We see the highly neutron-rich core at the center, the moderately neutron-rich disk in the equatorial plane, and higher $Y_{e}$ outflows in the polar regions. Figure reproduced from [40]
In the accretion disk, charged current reactions are initially fast enough for the matter to reach its equilibrium composition of $Y_{e}\sim(0.1-0.2)$ within a few milliseconds. At later times, the composition of the disk evolves because the equilibrium composition varies with the density and temperature of the matter, increasing with decreasing density. After $O(0.1\,{\rm s})$, weak interactions become too slow to maintain the equilibrium electron fraction, and the $Y_{e}$ of the disk freezes out. The expected value of $Y_{e}$ at freeze-out is $\sim 0.3$ [73, 74]. This is a particularly important process when considering r-process nucleosynthesis and kilonova emission. As mentioned earlier, $\sim 5\%-20\%$ of the initial accretion disk is unbound during the ADAF phase, and that matter will be ejected with an electron fraction close to the freeze-out value.
@@ Acknowledgements.
We thank Rodrigo Fernandez, Irene Tamborra, Somdutta Ghosh, Christian Fischer, and Damiano Forillo for discussions and comments provided on drafts of this chapter. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. FF gratefully acknowledges support from the Department of Energy, Office of Science, Office of Nuclear Physics, under contract number DE-AC02-05CH11231 and DE-SC0025023, from NASA through grant 80NSSC22K0719, and from the NSF through grant AST-2107932.